# SimGRAG: Leveraging Similar Subgraphs for Knowledge Graphs Driven Retrieval-Augmented Generation

### v1

@@ 6 Experiments
To evaluate the performance compared with existing approaches, we conduct extensive experiments on two tasks, i.e., Knowledge Graph Question Answering (KGQA) and Fact Verification.
@@ 6.1 Tasks and Datasets
Knowledge Graph Question Answering. We use the MoviE Text Audio QA dataset (MetaQA) Zhang et al. (2018) in this task, which provides a knowledge graph related to the field of movies. All the queries in the test set are adopted for evaluation, consisting of Vanilla 1-hop, 2-hop, and 3-hop question-answering in the same field.
Fact Verification. We adopt the FactKG dataset Kim et al. (2023b) in this task. It contains various statements including colloquial and written style claims, which can be verified using the DBpedia Lehmann et al. (2015). All statements in the test set are used in the evaluation, and a method should return Supported or Refuted after verification.
@@ 6.2 Baselines
Supervised task-specific models. State-of-the-art models for KGQA include EmbedKGQA Saxena et al. (2020), NSM He et al. (2021), and UniKGQA Jiang et al. (2022). They are trained on the MetaQA training set and evaluated by the test accuracy. For the task of fact verification, the KG version of GEAR Zhou et al. (2019) is included and trained on the FactKG training set.
Pre-trained LLMs. For both tasks, we evaluate two popular LLMs, ChatGPT OpenAI (2024) and Llama 3 70B Dubey et al. (2024), using 12-shots without any provided evidence.
KG-driven RAG with training. Recent method KELP Liu et al. (2024) trains the retriever over the training set, while G-retriever He et al. (2024) trains a graph neural network (GNN) to integrate query texts and subgraph evidences.
KG-driven RAG without training. Both KAPING Baek et al. (2023) and KG-GPT Kim et al. (2023a) only require retrieval subgraphs from the KGs without any training or fine-tuning.
Method MetaQA (Hits@1) FactKG
1-hop 2-hop 3-hop (Accuracy)
Supervised task-specific methods
EmbedKGQA 97.5 98.8 94.8 -
NSM 97.1 99.9 98.9 -
UniKGQA 97.5 99.0 99.1 -
GEAR - - - 77.7
Pre-trained LLMs
ChatGPT 60.0 23.0 38.7 68.5
Llama 3 70B 56.7 25.2 42.3 68.4
KG-driven RAG with training (Llama 3 70B)
KELP 94.7 96.0 - 73.3
G-Retriever 98.5 87.6 54.9 61.4
KG-driven RAG without training (Llama 3 70B)
KAPING 90.8 71.2 43.0 75.5
KG-GPT 93.6 93.6 88.2 69.5
SimGRAG (ours) 98.0 98.4 97.8 86.8
Table 1: Performance comparison of different approaches, where denotes the presence of entity leaks.
@@ 6.3 Comparative Results
As summarized in Table 1, supervised task-specific methods outperform KG-driven RAG approaches that require additional training, except on the MetaQA 1-hop dataset, where G-retriever achieves a 1% performance gain. Notably, supervised task-specific methods generally require smaller model sizes and lower training costs, making them a more cost-effective option in practice.
To avoid the costs of training, preparing samples, or serving a trained model, the inherent capabilities of pre-trained LLMs simplify the process, especially with the widely available APIs. However, directly using LLMs leads to the poorest performance. Thus, it is necessary to develop KG-driven RAG methods that do not require additional training. Among these methods, the proposed SimGRAG approach demonstrates substantially higher Hits@1 and accuracy in most cases. In fact, SimGRAG performs comparably to supervised task-specific models and even outperforms the supervised GEAR method on the FactKG dataset, while maintaining the plug-and-play usability.
Moreover, the performance gap between SimGRAG and other RAG approaches becomes larger as the complexity of the questions increases on the MetaQA dataset. As discussed in Section 4.2, the combined use of graph isomorphism and semantic similarity effectively reduces noise and ensures conciseness, thus benefiting the performance of SimGRAG for 2-hop and 3-hop questions. In contrast, KELP fails to support 3-hop paths due to the challenges of handling the vast number of potential paths with its trained path selection model.
@@ 6.4 Error Analysis
Table 2 summarizes the error distribution across the three main steps of the SimGRAG method. For detailed error examples, please refer to Appendix D.
The majority of errors occur during the query-to-pattern alignment step. In this stage, the large language model (LLM) frequently fails to follow the given instructions and examples, thereby generating the undesired pattern graphs. As the complexity of the queries increases in the MetaQA dataset, we also observe a higher incidence of errors in the subgraph-augmented generation step, since it is more difficult for the LLM to accurately extract relevant information for a complex question from the retrieved subgraphs. Also, errors are also encountered during the pattern-to-subgraph alignment phase on the FactKG dataset. In these cases, while the LLM generates reasonable subgraphs in line with the guidance, mismatches occur because the ground-truth subgraphs have different structures and thus cannot be successfully aligned, as illustrated in Appendix D.
6.5 Ablation Studies
MetaQA (Hits@1) FactKG
1-hop 2-hop 3-hop (Accuracy)
shot=4 98.6 96.5 92.8 84.0
shot=8 98.3 96.4 98.8 87.9
shot=12 98.0 98.4 97.8 86.8
$k=1$ 95.2 98.2 97.0 88.1
$k=2$ 98.0 97.9 97.6 87.6
$k=3$ 98.0 98.4 97.8 86.8
Table 3: Performance of the SimGRAG method by varying the number of few-shot examples and the parameter $k$ for semantic guided subgraph retrieval.
Few-shot in-context learning. Table 3 evaluates the proposed SimGRAG method by varying the number of examples in the prompts for the LLM, used in both pattern-to-graph alignment and verbalized subgraph-augmented generation. For the simplest MetaQA 1-hop questions, performance is not sensitive to the number of shots. In contrast, for more complex queries like those in the MetaQA 3-hop and FactKG datasets, we observe significant improvements when increasing from 4 to 8 shots.
Parameter $k$ for semantic guided subgraph retrieval. Table 3 reports the impact of parameter $k$ for retrieving top-$k$ subgraphs with the smallest graph semantic distance. For MetaQA 1-hop questions, setting $k=1$ leads to a significant drop in Hits@1, since many movies share the exactly same title, and retrieving fewer subgraphs makes it more difficult to cover the ground-truth answer. For MetaQA 2-hop and 3-hop questions, the choice of $k$ has a negligible impact on performance. Conversely, increasing $k$ leads to a slight decrease in accuracy on the FactKG dataset, since the top-1 subgraph is often sufficient and including more subgraphs will introduce noise for LLM.
@@ 6.5 Ablation Studies
Figure 5: Pareto optimal curves for retrieval.
@@ 6.6 Retrieval Efficiency
As discussed in Section 5, prior to running the optimized retrieval algorithm for top-$k$ subgraphs, we first perform a vector search to obtain the top-$k^{(n)}$ candidate nodes and top-$k^{(r)}$ candidate relations. Table 4 reports the average retrieval time per query, in which the vector search time dominates the total time. On the 10-million-scale DBpedia KG from the FactKG dataset, the overall retrieval time is 0.74 seconds per query, highlighting the efficiency and scalability of the optimized retrieval algorithm.
Additionally, we conduct a grid search over the parameters $k^{(n)}$ and $k^{(r)}$ to compare the top-$k$ retrieval and the optimized retrieval algorithms. Specifically, we randomly sample 100 queries that correctly generate patterns and manually identify the ground truth subgraphs for each query to evaluate retrieval performance using retrieval Hits@1. For detailed setups, please refer to Appendix E. By varying $k^{(n)}$ and $k^{(r)}$, Figure 5(a) presents the Pareto optimal curves, which plot the trade-off between average retrieval time and retrieval Hits@1. The results clearly show that the optimized retrieval algorithm significantly improves the performance, particularly in scenarios where a higher retrieval Hits@1 is desired in practice. Also, Figure 5(b) shows the overall latency for the proposed SimGRAG method, consisting of query-to-pattern alignment, pattern-to-subgraph alignment, and the verbalized subgraph-augmented generation. The top-$k$ retrieval algorithm without optimization dominates the overall latency when targeting higher retrieval Hits@1, while the optimized algorithm guarantees reasonable latency, surpassing even some fully-supervised models.
@@ 7 Conclusion
In this paper, we investigate the problem of KG-driven RAG and introduce a novel SimGRAG approach that effectively aligns query texts with KG structures. For query-to-pattern alignment, we employ an LLM to generate a pattern graph that aligns with the query text, ensuring plug-and-play usability without any training. For pattern-to-subgraph alignment, we introduce the Graph Semantic Distance (GSD) metric to quantify the alignment between the desired pattern and the underlying subgraphs in the KG, ensuring context conciseness and preventing entity leaks. Additionally, we propose an optimized algorithm to retrieve the top-$k$ similar subgraphs with the smallest GSD, improving retrieval efficiency and scalability. Extensive experiments demonstrate that SimGRAG consistently outperforms existing KG-driven RAG approaches.
8 Limitations
Since the proposed SimGRAG method is designed to offer plug-and-play usability, its performance is closely tied to the underlying capabilities of the large language model (LLM). Specifically, the method relies heavily on the ability of LLMs to understand and follow instructions effectively in both steps of the query-to-pattern alignment and verbalized subgraph-augmented generation. Thus, the performance of SimGRAG can be substantially degraded when utilizing lower-quality or less capable LLMs, especially in scenarios involving more complex queries that demand advanced reasoning skills.
@@ A.1 Knowledge Graph Question Answering
For the task of Knowledge Graph Question Answering, we use the MoviE Text Audio QA dataset (MetaQA), which is designed for research on question-answering systems on knowledge graphs Zhang et al. (2018). It provides a knowledge graph about movies, where entities include movie names, release years, directors, and so on, while the relations include starred actors, release year, written by, directed by, and so on. The queries are composed of Vanilla 1-hop, 2-hop, and 3-hop question answering in the field of movies. For the test set of MetaQA dataset, there are 9,947 questions for 1-hop, 14,872 for 2-hop, and 14,274 for 3-hop. Table 6 shows some example triples in the knowledge graph provided in the MetaQA Zhang et al. (2018) dataset, while Table 7 are some example questions in the dataset. The MetaQA dataset is released under the Creative Commons Public License.

### v2

@@ 6 Experiments
We conduct experiments on tasks of Knowledge Graph Question Answering (KGQA) and Fact Verification.
Method MetaQA (Hits@1) PQ (Hits@1) WC2014 FactKG
1-hop 2-hop 3-hop 2-hop 3-hop (Hits@1) (Accuracy)
Supervised task-specific methods
EmbedKGQA 97.5 98.8 94.8 - - - -
NSM 97.1 99.9 98.9 - - - -
UniKGQA 97.5 99.0 99.1 - - - -
Transfernet 97.5 100 100 - - - -
GEAR - - - - - - 77.7
Pre-trained LLMs
ChatGPT 60.0 23.0 38.7 - - - 68.5
Llama 3 70B 56.7 25.2 42.3 - - - 68.4
KG-driven RAG with training (Llama 3 70B)
KELP 94.7 96.0 - - - - 73.3
G-Retriever 98.5 87.6 54.9 61.8 46.7 67.5 61.4
KG-driven RAG without training (Llama 3 70B)
KAPING 90.8 71.2 43.0 41.0 52.1 88.1 75.5
KG-GPT 93.6 93.6 88.2 86.1 42.5 71.1 69.5
SimGRAG (ours) 98.0 98.4 97.8 88.7 78.6 98.1 86.8
Table 1: Performance comparison of different approaches, where denotes we provide oracle entities as it is the default setting of a method. Each reported value serves as an upper bound for the result obtained without oracle entities. Appendix D presents more discussions.
@@ 6.1 Tasks and Datasets
Knowledge Graph Question Answering.
We use the MoviE Text Audio QA dataset (MetaQA) Zhang et al. (2018) related to the field of movies. All the queries in the test set are adopted for evaluation, consisting of Vanilla 1-hop, 2-hop, and 3-hop question-answering in the same field. We also use the PathQuestions dataset (PQ) Zhou et al. (2018) developed from Freebase Bollacker et al. (2008) consisting of 2-hop and 3-hop queries, and the WorldCup2014 dataset (WC2014) Zhang et al. (2016) with sports-domain KGs.
Fact Verification.
We adopt the FactKG dataset Kim et al. (2023b), in which colloquial and written style claims can be verified using the DBpedia Lehmann et al. (2015). All statements in the test set are used in the evaluation, and a method should return Supported or Refuted after verification.
@@ 6.2 Baselines
Supervised task-specific models.
State-of-the-art models for KGQA include EmbedKGQA Saxena et al. (2020), NSM He et al. (2021), UniKGQA Jiang et al. (2022), and Transfernet Shi et al. (2021). They are trained on the MetaQA training set and evaluated by the test accuracy. For fact verification, the KG version of GEAR Zhou et al. (2019) is trained on the FactKG training set.
Pre-trained LLMs.
For both tasks, we evaluate two popular LLMs, ChatGPT OpenAI (2024) and Llama 3 70B Dubey et al. (2024), using 12-shots without any provided evidence.
KG-driven RAG with training.
Recent method KELP Liu et al. (2024) trains the retriever over the training set, while G-retriever He et al. (2024) trains a graph neural network (GNN) to integrate query texts and subgraph evidences.
KG-driven RAG without training.
Both KAPING Baek et al. (2023) and KG-GPT Kim et al. (2023a) only require retrieval subgraphs from the KGs without any training or fine-tuning.
MetaQA (Hits@1) PQ (Hits@1) WC2014 FactKG
1-hop 2-hop 3-hop 2-hop 3-hop (Hits@1) (Accuracy)
shot=4 98.6 96.5 92.8 90.9 78.3 97.2 84.0
shot=8 98.3 96.4 98.8 90.3 79.5 96.3 87.9
shot=12 98.0 98.4 97.8 88.7 78.6 98.1 86.8
$k=1$ 95.2 98.2 97.0 90.4 76.9 93.2 88.1
$k=2$ 98.0 97.9 97.6 90.3 77.7 93.6 87.6
$k=3$ 98.0 98.4 97.8 88.7 78.6 98.1 86.8
Llama3-70B 98.0 98.4 97.8 88.7 78.6 98.1 86.8
Phi4-14B 92.7 99.5 90.8 92.2 83.4 91.6 86.1
Qwen2.5-72B 98.6 99.8 98.2 88.7 77.7 97.5 83.6
Table 2: Performance of the SimGRAG method by varying the number of few-shot examples, the parameter $k$ for semantic guided subgraph retrieval, and different LLMs.
@@ 6.3 Comparative Results
As summarized in Table 1, supervised task-specific methods outperform KG-driven RAG approaches that require additional training. Notably, supervised task-specific methods generally require smaller model sizes and lower training costs, making them a more cost-effective option in practice.
Directly using LLMs leads to the poorest performance. As for KG-driven RAG methods without additional training, SimGRAG shows substantially higher Hits@1 and accuracy in most cases. In fact, SimGRAG performs comparably to supervised task-specific models and even outperforms the supervised GEAR method on the FactKG dataset.
Moreover, the performance gap between SimGRAG and other RAG approaches becomes larger as the complexity of the questions increases on the MetaQA dataset. As discussed in Section 4.2, the combined use of graph isomorphism and semantic similarity effectively reduces noise and ensures conciseness, thus benefiting the performance of SimGRAG for 2-hop and 3-hop questions.
6.4 Ablation Studies
Few-shot in-context learning.
Table 2 evaluates SimGRAG method by varying the number of examples in the prompts, used in both pattern-to-graph alignment and verbalized subgraph-augmented generation. For the simplest MetaQA 1-hop questions, performance is not sensitive to the number of shots. In contrast, for more complex queries like those in the MetaQA 3-hop, PQ 3-hop, and FactKG datasets, we observe significant improvements when increasing from 4 to 8 shots.
Parameter $k$ for semantic guided subgraph retrieval.
Table 2 reports the impact of parameter $k$ for retrieving top-$k$ subgraphs with the smallest graph semantic distance. For MetaQA 1-hop questions, setting $k=1$ leads to a significant drop in Hits@1, since many movies share exactly the same title, and retrieving fewer subgraphs makes it more difficult to cover the ground-truth answer. For MetaQA 2-hop and 3-hop questions, the choice of $k$ has a negligible impact on performance. Conversely, increasing $k$ leads to a slight decrease in accuracy on the FactKG dataset, since the top-1 subgraph is often sufficient and including more subgraphs will introduce noise for LLM.
Choice of Large Language Models.
We also evaluate the proposed SimGRAG using two additional open-source LLMs, including Phi4-14B Abdin et al. (2024) and Qwen2.5-72B Qwen et al. (2025). The results in Table 2 demonstrate that SimGRAG is generally robust across different LLMs. Even using the Phi-4 14B model without any training or finetuning, SimGRAG remains competitive with existing methods. Also, SimGRAG offers a plug-and-play solution on human-understandable KGs across various LLMs, and we expect its performance to improve with future LLM advancements.
Query pattern structure.
As outlined in Appendix F, we categorize query pattern structures into six classes and show the distributions in Table 3. Table 1 confirm that SimGRAG outperforms RAG baselines on multi-hop path queries, and it is also better on WC2014 dataset that contains 22% 2-hop conjunction queries. By further experiments on each category of queries for FactKG dataset, SimGRAG achieves the accuracies of 89%, 88%, and 85% on 2-hop conjunction, 3-hop conjunction, and 3-hop star queries, respectively.
Dataset Path Conjunction Star
1-hop 2-hop 3-hop 2-hop 3-hop 3-hop
MetaQA 1-hop 100% 0 0 0 0 0
MetaQA 2-hop 0 100% 0 0 0 0
MetaQA 3-hop 0 0 100% 0 0 0
PQ 2-hop 0 100% 0 0 0 0
PQ 3-hop 0 0 100% 0 0 0
WC2014 64% 14% 0 22% 0 0
FactKG 32% 28% 5% 17% 8% 10%
Table 3: Distribution of query pattern structures.
@@ 6.5 Error Analysis
Table 4 summarizes the error distribution across the three main steps of the SimGRAG method. For detailed error examples, please refer to Appendix E.
Many errors occur during the query-to-pattern alignment step, where the LLM fails to follow the given instructions and examples, thereby generating the undesired pattern graphs. Generally, both 2-hop and 3-hop queries roughly have consistent proportion of errors. But there are more errors on 1-hop queries, since we use the same few-shot examples for all MetaQA queries, which are all about 2-hop or 3-hop patterns. They make the LLM sometimes generate 2-hop patterns for 1-hop queries.
As the complexity of the queries increases in the MetaQA dataset, we also observe a higher incidence of errors in the subgraph-augmented generation step, since it is more difficult for the LLM to accurately extract relevant information for a complex question from the retrieved subgraphs.
On the FactKG dataset, errors are also encountered during the pattern-to-subgraph alignment. In these cases, while the LLM generates reasonable subgraphs in line with the guidance, mismatches occur because the ground-truth subgraphs have different structures and thus cannot be successfully aligned, as illustrated in Appendix E.
@@
Figure 5: Pareto optimal curves for retrieval.
@@
Method MetaQA FactKG
1-hop 2-hop 3-hop
KELP 3.6 4.8 - 5.4
G-Retriever 4.1 4.3 4.4 5.4
KAPING 3.3 5.9 8.8 10.6
KG-GPT 10.1 12.3 13.1 13.3
Ours (4-shot) 5.5 5.9 7.3 10.2
Ours (12-shot) 9.0 9.1 11.9 14.2
Table 6: Comparison of average query latency (s).
@@ 6.6 Retrieval Efficiency
As discussed in Section 5, we first perform a vector search to obtain the top-$k^{(n)}$ candidate nodes and top-$k^{(r)}$ candidate relations. Table 5 reports the average retrieval time per query, in which the vector search time dominates the total time. On the 10-million-scale DBpedia KG from the FactKG dataset, the overall retrieval time is 0.74 seconds per query, highlighting the efficiency and scalability of the optimized retrieval algorithm.
Additionally, we conduct a grid search over the parameters $k^{(n)}$ and $k^{(r)}$ to compare the top-$k$ retrieval and the optimized algorithms. Please refer to Appendix G for detailed setups. Figure 5(a) presents the Pareto optimal curves, which plot the trade-off between average retrieval time and retrieval Hits@1. The results clearly show that the optimized retrieval algorithm significantly improves the performance, particularly in scenarios where a higher retrieval Hits@1 is desired in practice. Also, Figure 5(b) shows the overall latency for the proposed SimGRAG method, in which the optimized algorithm guarantees reasonable latency.
6.7 Overall Latency
We run each method on a NVIDIA A6000 GPU using Ollama 4-bit quantization for Llama3 70B. Table 6 reports the average latency for answering each query. Generally, our method has similar latency compared with others using the default 12-shot in-context learning. It could be much faster with 4-shot learning while still providing competitive performance, as confirmed by Table 2.
@@ 7 Conclusion
In this paper, we investigate the problem of KG-driven RAG and introduce a novel SimGRAG approach that effectively aligns query texts with KG structures. For query-to-pattern alignment, we employ an LLM to generate a pattern graph that aligns with the query text. For pattern-to-subgraph alignment, we introduce the Graph Semantic Distance (GSD) metric to quantify the alignment between the desired pattern and the underlying subgraphs in the KG. Additionally, we propose an optimized algorithm to retrieve the top-$k$ similar subgraphs with the smallest GSD, improving retrieval efficiency and scalability. Extensive experiments demonstrate that SimGRAG consistently outperforms existing KG-driven RAG approaches.
Acknowledgments
This work was substantially supported Key Projects of the National Natural Science Foundation of China (Grant No. U23A20496). Weiguo Zheng is the corresponding author.
Limitations
The performance of SimGRAG method is closely tied to the underlying capabilities of the large language model (LLM). Specifically, the method relies heavily on the ability of LLMs to understand and follow instructions effectively in both steps of the query-to-pattern alignment and verbalized subgraph-augmented generation. Thus, the performance of SimGRAG can be substantially degraded when utilizing lower-quality or less capable LLMs, especially in scenarios involving more complex queries that demand advanced reasoning skills.
Furthermore, following the characteristics of KGs used by existing studies Baek et al. (2023); Kim et al. (2023a); Liu et al. (2024); He et al. (2024), we also assume that our input KG aligns with human cognition. It is a key requirement for the plug-and-play usability for the SimGRAG method. However, when using industrial domain-specific KGs which diverge significantly from commonly used schemas, it is challenging for LLMs to predict the desired nodes, edges or pattern structures during the query-to-pattern alignment stage. It is still under exploration how effectively fine-tuning LLMs can help to align generated patterns with such special KG structures. Also, we can include the specialized KG schema in prompts, guiding LLMs to generate patterns more likely isomorphic to desired subgraphs in the KG.
Additionally, for domain-specific KGs, linking query entities to corresponding candidate entities in the KG could be challenging, particularly when the embedding model has not been trained on such data. Therefore, rather than relying on a plug-and-play embedding model, future work may fine-tune the embedding models on domain-specific data or explore alternative entity linking approaches.
@@ A.1 Knowledge Graph Question Answering
For the task of Knowledge Graph Question Answering, we use the MoviE Text Audio QA dataset (MetaQA) Zhang et al. (2018), PathQuestions dataset (PQ) Zhou et al. (2018), and the WorldCup2014 dataset (WC2014) Zhang et al. (2016).
MoviE Text Audio QA dataset (MetaQA) is designed for research on question-answering systems on knowledge graphs Zhang et al. (2018). It provides a knowledge graph about movies, where entities include movie names, release years, directors, and so on, while the relations include starred actors, release year, written by, directed by, and so on. The queries are composed of Vanilla 1-hop, 2-hop, and 3-hop question answering in the field of movies. For the test set of MetaQA dataset, there are 9,947 questions for 1-hop, 14,872 for 2-hop, and 14,274 for 3-hop. Table 8 shows some example triples in the knowledge graph provided in the MetaQA Zhang et al. (2018) dataset, while Table 9 are some example questions in the dataset. The MetaQA dataset is released under the Creative Commons Public License.
PathQuestions dataset (PQ) is built on Freebase KG Bollacker et al. (2008), which contains 1,908 2-hop path queries and 5,198 3-hop path queries Zhou et al. (2018). Table 10 shows some example triples in the PathQuestions knowledge graph, while Table 11 are some example questions. It is under a Creative Commons Attribution 4.0 International Licence.
WorldCup2014 dataset (WC2014) contains a knowledge graph about football players that participated in FIFA World Cup 2014 Zhang et al. (2016). There are 10,162 queries of WC2014, which is a mixture of 6,482 single-relation questions, 1,472 two-hop path questions, and 2,208 conjunctive questions. Table 12 shows some example triples in the WC2014 knowledge graph, while Table 13 are some example questions. It is under a Creative Commons Attribution 4.0 International Licence.
     Head      Relation      Tail
     eleanor_of_provence      children      beatrice_of_england
     manuel_i_of_portugal      gender      male
     joan_crawford      spouse      phillip_terry
     barbara_of_portugal      spouse      ferdinand_vi_of_spain
     empress_myeongseong      cause_of_death      regicide
     frederica_of_mecklenburg-strelitz      spouse      ernest_augustus_i_of_hanover
     henri_victor_regnault      gender      male
     adelaide_of_lowenstein_wertheim_rosenberg      children      maria_josepha_of_portugal
     prince_frederick_duke_of_york_and_albany      place_of_death      london
     mary_boleyn      spouse      william_carey_1490
Table 10: Example triples in the knowledge graph of PQ dataset.
Dataset Example questions
PQ 2-hop 1. john_b_kelly_sr’s son’s job?
2. what is the sex of spouse of mary_stuart_countess_of_bute?
3. where does virginia_heinlein’s spouse work for?
PQ 3-hop 1. who is the mom of father of mary_of_teck’s heir?
2. what is the name of the gender of son of henry_i_of_england’s mother?
3. ferdinand_ii_of_aragon’s parent’s heir’s nation?
Table 11: Example questions in the PQ dataset.
            Head             Relation             Tail
            Esseid_BELKALEM             plays_in_club             Watford_FC
            Frank_LAMPARD             plays_for_country             England
            JOAO_MOUTINHO             plays_in_club             AS_Monaco
            Agustin_ORION             plays_position             Goalkeeper
            Rickie_LAMBERT             is_aged             32
            Pedro_RODRIGUEZ             plays_position             Forward
            HENRIQUE             is_aged             27
            OGC_Nice             is_in_country             France
            Andres_GUARDADO             wears_number             18
            Michel_VORM             plays_position             Goalkeeper
Table 12: Example triples in the knowledge graph of WC2014 dataset.
Dataset Example questions
WC2014 1-hop 1. which football club does Alan_PULIDO play for?
2. what position does Alan_PULIDO play?
3. which country is the soccer team Atletico_Madrid based in?
WC2014 2-hop 1. which professional foootball team has a player from Belgium?
2. where is the football club that Rafael_MARQUEZ plays for?
3. which country does Mathieu_DEBUCHY play professional in?
WC2014 Conjunction 1. name a player who plays at Forward position at the club Tigres_UANL?
2. who are the Italy players at club US_Citta_di_Palermo?
3. which Portugal footballer plays at position Goalkeeper?
Table 13: Example questions in the WC2014 dataset.